### v1

@@ Abstract
This study provides a comprehensive investigation into the fracture behavior of irradiated Fe–Ni–Cr alloys, focusing on the radiation-induced ductile-to-brittle transition (DBT). Unlike the commonly studied temperature-driven DBT, this transition is driven by radiation-induced defects, which impede dislocation motion and reduce the material’s capacity for plastic deformation, ultimately leading to brittle failure. Utilizing Molecular Dynamics simulations, this work examines the interaction between radiation-induced defects, such as dislocation loops and voids, and their influence on crack propagation and energy dissipation mechanisms. The results reveal distinct roles played by these defects: voids facilitate crack growth by reducing local cohesive energy, while dislocation loops act as barriers that impede crack advancement and redirect crack paths, significantly altering crack morphology. Building on the classical approach of separating fracture energy into brittle cleavage and plastic components, this study adapts this decomposition to irradiated materials. This approach appraises the distinct roles of surface energy and plastic work across varying radiation damage levels, providing critical insights into how irradiation-induced defects govern fracture dynamics.
@@ 1 Introduction
The embrittlement of metals, resulting from reduced plastic deformation preceding fracture, is a significant determinant of the durability and safety of structural materials, particularly under the extreme conditions encountered in nuclear reactors. Irradiation embrittlement plays a crucial role in the degradation of the mechanical properties of these materials. Radiation defects arise when energetic particles, such as neutrons or ions, collide with atoms in the lattice, displacing them from their original positions and creating vacancies and interstitial atoms, which subsequently cluster into more complex defect structures such as dislocations and voids [1, 2, 3, 4].
In irradiated materials, ductile-to-brittle transition (DBT) is driven by radiation defects, which impede dislocation motion and reduce the material’s capacity for plastic deformation, ultimately leading to a brittle failure.
@@
DBT is a well-known phenomenon in metals that is usually related to the temperature sensitivity of fracture stress and yield stress [5, 6, 7, 8]. DBT is characterized by a shift from ductile fracture, which involves significant plastic deformation, to brittle fracture, where materials fail with little prior deformation [9, 10]. This transition is heavily influenced by temperature, stress state, and microstructural factors, particularly the presence of defects. While temperature is recognized as an important factor in the overall understanding of DBT phenomenon, the accumulation of radiation-induced defects and their influence on mechanical properties are considered the determining elements in the case of irradiated materials.
@@
Fig. 1 illustrates the impact of irradiation on the mechanical behavior of stainless steel. The stress–strain curves demonstrate the reduction of ductility with increasing radiation-induced porosity (compare Fig. 1a). The curves clearly show a decrease in elongation at fracture with higher radiation doses, indicating a significant loss of plasticity. This loss of ductility is directly related to the increase of radiation porosity and defect accumulation.
@@
Figure 1: a) Engineering stress–strain curves of stainless steel after irradiation, demonstrating ductile-to-brittle transition (after Byun et al. [11]). b) Transmission electron microscopy (TEM) analysis of steel irradiated to 0.12 dpa, showing radiation-induced porosity in the form of defect clusters at varying magnifications [12]. c) TEM images illustrating the dislocation-blocking mechanism, where a dislocation line encircles radiation defects at varying magnifications [12]. These images depict the irradiated material after deformation, highlighting the interaction between dislocations and radiation-induced defects.
@@
The reduction of ductility is a critical factor contributing to ductile–brittle transition (DBT). As shown in Fig. 1b), irradiation-induced defects such as voids and dislocation loops form clusters that disrupt the crystalline structure, impeding the material’s ability to deform plastically [13, 14, 12]. Additionally, TEM images in Fig. 1c) reveal the dislocation-blocking mechanism, where dislocation lines bend and encircle radiation defects, forming curved arms. This interaction, observed after deformation, significantly restricts plastic flow by impeding dislocation motion [15, 16, 17, 18]. These microstructural changes, driven by the accumulation of defects, significantly degrade the mechanical integrity of the material under applied stress. Thus, defect accumulation can play a pivotal role in driving DBT.
@@
Molecular dynamics (MD) simulations play a critical role in studying DBT by providing atomic-level insights into the mechanisms driving embrittlement [19, 20]. In the case of irradiated materials, MD allows for detailed investigation of radiation-induced defects and their interactions with crack propagation and dislocation motion [15, 21, 22, 17, 18, 14, 23]. These simulations are particularly useful for examining the effects of radiation defects on material behavior, enabling the exploration of how defect accumulation changes the mechanical properties of the materials [16, 17]. MD can simulate extreme irradiation conditions and provide a clear understanding of the transition from ductile to brittle fracture at various defect densities.
@@
This paper presents a novel investigation into the role of radiation-induced defects, expressed by the displacement per atom (dpa) parameter [4], in driving the ductile-to-brittle transition. Unlike previous studies, which primarily focus on temperature-driven DBT, this work examines the direct influence of irradiation-induced defects on dislocation dynamics and energy dissipation mechanisms during crack propagation. By systematically analyzing how varying dpa levels govern defect accumulation and alter the material’s capacity for plastic deformation, this study establishes a direct link between irradiation damage and fracture behavior, an aspect that has not been comprehensively addressed before. Furthermore, this paper provides the first detailed correlation between defect density and the microstructural mechanisms governing DBT. Through a combination of MD simulations and advanced defect characterization techniques, this study offers original insights into how variations in dpa influence the mechanical response, including the transition from ductile to brittle fracture.
@@
A comprehensive understanding of DBT phenomenon in irradiated materials is crucial for accurately predicting material performance in nuclear applications and for devising effective strategies to mitigate embrittlement effects. Understanding the fracture behavior of irradiated structural materials is essential to ensuring the safety and reliability of structural components of nuclear reactors operating in exceptionally harsh environments due to the combination of high stresses, a chemically aggressive coolant, high or, in the case of the accelerator system, extremely low temperatures, and intense radiation fluxes [24, 25, 26, 22, 27].
@@ 2.1 Preparation of the irradiated samples
Figure 2: Structural evolution of radiation defects in Fe-Ni-Cr Alloy with increasing dpa. The legend identifies the types of dislocation loops and their designations. Voids are represented in blue.
Figure 3: Dislocation loops (a) and voids density (b) as a function of the irradiation dose in dpa obtained from MD simulations.
For this study, a Cr-rich fcc $\rm Fe_{55}Ni_{19}Cr_{26}$ sample with dimensions of $137.5\times 141.0\times 148.0$$\text{\AA{}}^{3}$ was prepared containing 262 080 atoms. The composition $\rm Fe_{55}Ni_{19}Cr_{26}$ represents the atomic percentage of each element, and the sample was oriented along the $\langle 001\rangle$ crystallographic direction. The irradiation of samples for a wide range of dpa was achieved through overlapping collision cascade simulations [3, 4, 12]. Each cascade was initiated by a primary knock-on atom (PKA) with an energy of 10 keV, randomly selected within the simulation box to ensure uniform damage distribution.
For example, to achieve dpa values of 0.008, 0.038, 0.152, and 0.266, as shown in Fig. 2; 20, 100, 400, and 700 consecutive recoil MD simulations were performed, respectively. Each recoil event was simulated for 20 ps under the NPT ensemble, followed by an additional 10 ps to allow for stress relaxation and swelling [12]. Defect formation, including dislocation nucleation and void generation, was analyzed across various dpa levels using the Dislocation Extraction Algorithm (DXA) in OVITO [36]. Dpa dose was calculated using the model proposed by Nordlund et al. [4].
The evolution of the density of radiation-induced dislocation loops and voids, as functions of the irradiation dose, are shown in Fig. 3a) and Fig. 3b), respectively. The density of radiation defects increases gradually at lower dpa, reflecting the early stages of dislocation loop nucleation and the initial aggregation of vacancies into stable void structures. At higher dpa levels, the accumulation and interaction of dislocations lead to the formation of various dislocation types and a more complex defect network. Fig. 2 provides a detailed depiction of how microstructural features, including dislocation loops and void densities, evolve under increasing radiation exposure, highlighting the transition from isolated defects to a dense defect structure.
@@ 2.2 Simulations of fracture
Figure 4: Schematic representation of the 3D atomic-scale fracture model. The simulation box prepared by replicating an irradiated sample along the x- and y-axes is shown on the left side. The model with an initial crack introduced at the midplane subjected to uniform elongation is shown on the right side.
@@
A three-dimensional (3D) atomic-scale model shown in Fig. 4 was prepared to enable the simulation and analysis of fracture mechanics phenomena. The irradiated samples obtained from the irradiation step (Section 2.1) were duplicated using periodic replication techniques along the x- and y- axes to achieve the desired model dimensions of $275\times 282\times 148$$\text{\AA}^{3}$. The simulation box, containing 1 048 320 atoms, provides a sufficient volume to capture crack-related phenomena without significant boundary interference. To ensure the stability of the fracture model under the specified conditions, a relaxation step was performed for 10 ps following an equilibration process using a canonical NPT ensemble for 100 ps at a target temperature of 300 K and zero stress. This equilibration phase allowed both atomic positions and simulation box dimensions to adjust dynamically, ensuring the system reached a stable thermodynamic state. Following equilibration, the NVT ensemble was applied to preserve a constant temperature throughout the deformation process, ensuring thermal stability and eliminating artifacts such as unintended energy dissipation or volume fluctuations that might influence crack propagation behavior. An initial crack with a length of 28 Å was then introduced at the midplane of the specimen by selectively removing the bonds between atoms along the x-axis to create a sharp discontinuity.
pristine sample
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
sample irradiated to 0.152 dpa
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 5: Crack propagation for non-irradiated and irradiated sample.
unirradiated sample
sample irradiated to 0.152 dpa
Figure 6: Equivalent shear strain fields calculated using the Green–Lagrangian strain tensor [37] in a selected cross-section of unirradiated and irradiated samples.
@@
Due to the low thickness of the crystal compared to its height and width, this problem is modeled as a constrained 3D system [2]. Periodic boundary conditions are applied along the thickness of the crystal (i.e., z-direction), while non-periodic boundary conditions are used in the x- and y-directions, corresponding to the crack propagation and loading direction, respectively. To prevent necking and minimize residual effects (unwanted structural deformations), the atoms at the left and right edges of the sample (see Fig. 4) are restricted from moving in the x-direction but remain free to move along the y- and z-axes, in accordance with the assumed potential and loading conditions. In addition, two atomic layers at the top and bottom of the sample along the y-direction are kept frozen. The system is subjected to uniform elongation along the y-direction, applied to the upper and lower boundaries. The displacement rate is set to 0.08 Å/ps, corresponding to a global strain rate of $10^{9}$ 1/s. Several alternative boundary condition configurations, such as fully periodic (ppp) [38] or fully free (ffp) and non-periodic with shrink-wrapped (ssp) [2, 20, 39, 6] systems, were also tested. However, in the case of (ffp) and (ssp) conditions, excessive necking of the sample occurred due to its high ductility, severely distorting the plastic zone and obscuring the underlying fracture mechanisms. On the other hand, (ppp) configurations artificially modify plastic flow near the crack front due to the interaction of dislocations with periodic boundary conditions, which create virtual copies of the sample that influence the dislocation fields. Therefore, the constrained boundary setup used in this work was selected as a compromise that preserves realistic plasticity and dislocation behavior while preventing unphysical deformation modes. Following the criteria established by Buehler [40], a time step of 1 fs is used for the simulations. This approach provides the foundation for analyzing the mechanical response and fracture behavior of both unirradiated and irradiated samples.
@@
The crack propagation in both a unirradiated sample and in a selected irradiated sample (0.152 dpa) is illustrated in Fig. 5, where the role of radiation-induced defects on dislocation behavior and crack path characteristics is highlighted. In the pristine sample, crack growth generates significant dislocation activity near the crack tip, dissipating energy through plastic deformation and effectively blunting the crack tip.
However, in the irradiated sample, the presence of radiation-induced defects modifies dislocation dynamics, disrupting the characteristic mechanisms of energy dissipation through plastic deformation. Fig. 6 presents shear strain fields in both unirradiated and irradiated samples across various time steps. Using the Green–Lagrangian strain tensor, the local shear strain is computed, and these scalar quantities are output as particle properties [41, 37]. The images show clear differences in strain localization patterns, where the irradiated samples exhibit concentrated strain fields at radiation-induced defects.
@@ 3 Characterization of mechanisms governing crack propagation in irradiated materials
In irradiated materials, crack propagation is profoundly influenced by the density, type, and spatial distribution of radiation-induced defects. These defects act as obstacles to dislocation motion, modifying the local stress fields around the crack and restricting the material’s ability to dissipate mechanical energy through plastic deformation. As a result, the crack path can be redirected, with propagation either hindered or accelerated depending on the nature and arrangement of the defects. To investigate these effects, the crack evolution at different irradiation levels is analyzed, focusing on crack morphology, and interactions with radiation-induced defects.
In the following subsections, various modes of crack propagation in irradiated materials are examined in detail. Subsection 3.1 examines characteristic modes of crack growth, highlighting how radiation-induced defects alter crack trajectories and lead to phenomena such as crack branching, void coalescence, and secondary crack formation. Subsection 3.2 delves into the fundamental mechanisms governing defect–crack interactions, focusing on how specific defects (e.g., voids, dislocation loops, rigid inclusions) either facilitate or impede crack propagation and how these interactions influence fracture behavior.

### v2

@@ Abstract
Neutron irradiation in structural alloys promotes defect clustering, which suppresses plasticity and triggers a ductile-to-brittle transition (DBT), a key degradation mechanism limiting fracture resistance in nuclear materials. This study investigates the fracture mechanisms underlying this transition in irradiated Fe–Ni–Cr alloys. Using Molecular Dynamics simulations, we examine how different defect types influence crack propagation and energy dissipation mechanisms. The results reveal distinct roles of these defects: voids facilitate crack growth by reducing local cohesive energy, while dislocation loops act as barriers that impede crack advancement and redirect crack paths, significantly altering crack morphology. Building on the classical approach of separating fracture energy into brittle cleavage and plastic components, this study adapts the decomposition to irradiated materials. This framework quantifies the evolving contributions of surface energy and plastic work across increasing radiation damage levels, providing critical insight into how irradiation-induced defects govern fracture dynamics.
@@ 1 Introduction
The embrittlement of metals, resulting from reduced plastic deformation preceding fracture, is a significant determinant of the durability and safety of structural materials, particularly under the extreme conditions encountered in nuclear reactors. Irradiation embrittlement plays a crucial role in the degradation of the mechanical properties of these materials [1, 2, 3]. Radiation defects arise when energetic particles, such as neutrons or ions, collide with atoms in the lattice, displacing them from their original positions and creating vacancies and interstitial atoms, which subsequently cluster into more complex defect structures such as dislocations and voids [4, 5, 6]. Those defects impede dislocation motion and reduce the material’s capacity for plastic deformation, ultimately leading to a brittle failure. The related phenomena are collectively referred to as the radiation-induced ductile-to-brittle transition (DBT).
@@
DBT is a well-known phenomenon in metals that is usually related to the temperature sensitivity of fracture stress and yield stress [7, 8, 9, 10]. DBT is characterized by a shift from ductile fracture, which involves significant plastic deformation, to brittle fracture, where materials fail with little prior deformation [11, 12]. This transition is heavily influenced by temperature, stress state, and microstructural factors, particularly the presence of defects. While temperature is recognized as an important factor in the overall understanding of the DBT phenomenon, the accumulation of radiation-induced defects and their influence on mechanical properties are considered the determining elements in the case of irradiated materials.
@@
Fig. 1 illustrates the impact of irradiation on the mechanical behavior of a stainless steel. The stress–strain curves demonstrate the reduction of ductility with increasing radiation-induced porosity (compare Fig. 1(a). The curves clearly show a decrease in elongation at fracture with higher radiation doses, indicating a significant loss of plasticity. This loss of ductility is directly related to the increase of radiation porosity and defect accumulation.
@@
Figure 1: (a) Engineering stress–strain curves of stainless steel after irradiation, demonstrating ductile-to-brittle transition (after Byun et al. [13]). (b) Transmission electron microscopy (TEM) analysis of steel irradiated to 0.12 dpa, showing radiation-induced porosity in the form of defect clusters at varying magnifications [14]. (c) TEM images illustrating the dislocation-blocking mechanism, where a dislocation line encircles radiation defects at varying magnifications [14]. These images depict the irradiated material after deformation, highlighting the interaction between dislocations and radiation-induced defects.
@@
The reduction of ductility is a critical factor contributing to ductile–brittle transition (DBT). As shown in Fig. 1(b), irradiation-induced defects such as voids and dislocation loops form clusters that disrupt the crystalline structure, impeding the material’s ability to deform plastically [15, 16, 14]. Additionally, TEM images in Fig. 1(c) reveal the dislocation-blocking mechanism, where dislocation lines bend and encircle radiation defects, forming curved arms. This interaction, observed after deformation, significantly restricts plastic flow by impeding dislocation motion [17, 18, 19]. These microstructural changes, driven by the accumulation of defects, significantly degrade the mechanical integrity of the material under applied stress. Thus, defect accumulation can play a pivotal role in driving DBT.
@@
Molecular dynamics (MD) simulations, with well-known limitations concerning spatial and temporal scales, play a critical role in studying DBT by providing atomic-level insights into the mechanisms driving embrittlement [20, 21, 22]. In the case of irradiated materials, MD enables in-depth analysis of how radiation-induced defects influence crack propagation pathways and fracture behavior [23, 24, 16, 25]. It further provides insight into the mechanisms by which these defects interact with dislocations and alter their mobility [26, 19, 27]. These simulations are particularly useful for examining the effects of radiation defects on material behavior, enabling the exploration of how defect accumulation changes the mechanical properties of the materials [18, 28, 19]. MD can simulate extreme irradiation conditions and provide a clear understanding of the transition from ductile to brittle fracture at various defect densities. Furthermore, MD results are essential inputs for multiscale models, enabling the development of constitutive laws that accurately represent defect-driven plasticity and fracture. That integration bridges atomic-scale mechanisms with continuum descriptions, improving predictions of damage evolution and DBT [29, 30, 31].
@@
This paper presents a novel investigation into the role of radiation-induced defects, expressed by the displacement per atom (dpa) parameter [1], in driving DBT. Unlike previous studies, which primarily focus on temperature-driven DBT, this work examines the direct influence of irradiation-induced defects on dislocation dynamics and energy dissipation mechanisms during crack propagation. By systematically analyzing how varying dpa levels govern defect accumulation and alter the material’s capacity for plastic deformation, this study establishes a direct link between irradiation damage and fracture behavior, an aspect that has not been comprehensively addressed before. Furthermore, this paper provides the first detailed correlation between defect density and the microstructural mechanisms governing DBT. Through a combination of MD simulations and advanced defect characterization techniques, this study offers original insights into how variations in dpa influence the mechanical response, including the transition from ductile to brittle fracture.
@@
A comprehensive understanding of the DBT phenomenon in irradiated materials is crucial for accurately predicting material performance in nuclear applications and for devising effective strategies to mitigate embrittlement effects. Understanding the fracture behavior of irradiated structural materials is essential to ensuring the safety and reliability of structural components of nuclear reactors operating in exceptionally harsh environments due to the combination of high stresses, a chemically aggressive coolant, high or, in the case of accelerator systems, extremely low temperatures, and intense radiation fluxes [32, 33, 34, 35].
The relevance of our results is particularly strong for structural materials in nuclear reactors, where neutron irradiation leads to volumetric defect accumulation throughout the bulk of the material. While the types of defects considered in this study can form under both neutron and ion irradiation, the key distinction lies in the spatial extent of damage. Our simulations are designed to reproduce the uniform, bulk-like defect distribution characteristic of neutron exposure, rather than the gradient defect profiles characteristic of layers in ion-irradiated samples.
@@ 2 Simulation methodology
This section outlines the molecular dynamics simulation approach used to investigate the effects of irradiation-induced defects on fracture behavior in Fe–Ni–Cr alloys. It is divided into two parts: first, the preparation of irradiated samples with controlled displacement damage levels is described; second, the methodology for simulating crack propagation in these pre-damaged samples is presented.
@@ 2.1 Preparation of the irradiated samples
To initiate the displacement cascade simulations, primary knock-on atoms (PKAs) were randomly selected from the simulation cell. Although the atomic masses of Fe, Cr, and Ni are relatively close, the initial velocity of each PKA was adjusted according to its mass to ensure a constant recoil energy of 10 keV, in accordance with the kinetic energy dependence on atomic mass. This approach allowed for a physically consistent treatment of recoil events across different atomic species. To achieve dpa values of 0.008, 0.038, 0.152, and 0.266, as shown in Fig. 2; 20, 100, 400, and 700 consecutive recoil MD simulations were performed, respectively. Each recoil event was simulated for 20 ps under the NPT ensemble, followed by an additional 10 ps to allow for stress relaxation and swelling [14]. Defect formation, including dislocation nucleation and void generation, was analyzed across various dpa levels using the Dislocation Extraction Algorithm (DXA) for dislocation analysis and Voronoi-based tessellation methods for identifying voids, as implemented in OVITO [44]. Dpa dose was calculated following the approach proposed by Nordlund et al. [1, Eq. (10)], using the expression $\text{dpa}=\frac{1}{N}\cdot\frac{0.8T_{\rm PKA}}{2E_{d}}$, where $T_{\rm PKA}$ denotes the kinetic energy of PKA and $N$ is the total number of atoms in the simulation cell. In this study, a constant value of displacement threshold energy $E_{\rm d}$ = 40 eV was used, which lies within the typical range reported for Fe, Ni, and Cr [45, 46]. While the model can accommodate element-specific $E_{\rm d}$ values, this approximation was adopted for consistency across the alloy system.
Fig. 2 illustrates how microstructural features, including dislocation loops and void densities, evolve under increasing radiation exposure, highlighting the transition from isolated defects to a dense defect structure. Additional details on the development of radiation-induced defect structures in the samples are provided in Appendix A.
Figure 2: Structural evolution of radiation defects in Fe–Ni–Cr alloy with increasing dpa. The legend identifies the types of dislocation loops and their designations. Voids are represented in burgundy.
@@ 2.2 Simulations of fracture
A three-dimensional (3D) atomic-scale model shown in Fig. 3 was prepared to enable the simulation and analysis of fracture mechanics phenomena. The irradiated samples obtained from the irradiation step (Section 2.1) were duplicated using periodic replication techniques along the $x$- and $y$-axes to achieve the desired model dimensions of $275\times 282\times 148$$\text{\AA}^{3}$. The simulation box, containing 1 048 320 atoms, provides a sufficient volume to capture crack-related phenomena without significant boundary interference. To ensure the stability of the fracture model under the specified conditions, a relaxation step was performed for 10 ps following an equilibration process using a canonical NPT ensemble for 100 ps at a target temperature of 300 K and zero stress. This equilibration phase allowed both atomic positions and simulation box dimensions to adjust dynamically, ensuring the system reached a stable thermodynamic state. Following equilibration, the NVT ensemble was applied to preserve a constant temperature throughout the deformation process, ensuring thermal stability and eliminating artifacts such as unintended energy dissipation or volume fluctuations that might influence crack propagation behavior. An initial crack with a length of 28 Å was introduced at the midplane of the specimen by selectively removing bonds between atoms along the $x$-axis over the full thickness of the specimen (i.e., across all atomic layers in the $z$-direction), creating a sharp discontinuity.
Figure 3: Schematic representation of the 3D atomic-scale fracture model. (a) The simulation box prepared by replicating an irradiated sample along the $x$- and $y$-axes. (b) The model with an initial crack introduced at the midplane subjected to uniform elongation.
Figure 4: Crack propagation for non-irradiated (top) and irradiated (bottom) sample.
@@
Due to the low thickness of the crystal compared to its height and width, this problem is modeled as a constrained 3D system [23]. Periodic boundary conditions are applied along the thickness of the crystal (i.e., $z$-direction), while non-periodic boundary conditions are used in the $x$- and $y$-directions, corresponding to the crack propagation and loading direction, respectively. To prevent necking and minimize residual effects (unwanted structural deformations), the atoms at the left and right edges of the sample (see Fig. 3) are restricted from moving in the $x$-direction but remain free to move along the $y$- and $z$-axes, in accordance with the assumed potential and loading conditions. In addition, two atomic layers at the top and bottom of the sample along the $y$-direction are kept frozen. The system is subjected to uniform elongation along the $y$-direction, applied to the upper and lower boundaries. The displacement rate is set to 0.08 Å/ps, corresponding to a global strain rate of $10^{9}$ 1/s. Following the criteria established by Buehler [47], a time step of 1 fs is used for the simulations. This approach provides the foundation for analyzing the mechanical response and fracture behavior of both unirradiated and irradiated samples.
A detailed analysis of alternative boundary condition configurations—fully periodic (ppp), fully free (ffp), and non-periodic shrink-wrapped (ssp)—is presented in Appendix B. This analysis evaluates their effect on plastic deformation and fracture behavior and provides the rationale for employing the constrained ssp setup in the main study.
@@
The crack propagation in both a unirradiated sample and in a selected irradiated sample (0.152 dpa) is illustrated in Fig. 4, where the role of radiation-induced defects on dislocation behavior and crack path characteristics is highlighted. In the pristine sample, crack growth generates significant dislocation activity near the crack tip, dissipating energy through plastic deformation and effectively blunting the crack tip. In the irradiated sample, radiation defects additionally increase the overall dislocation density in the whole volume. Strain-field data further illustrating the differences between the deformation responses of unirradiated and irradiated materials is presented in Appendix C.
@@ 3 Characterization of mechanisms governing crack propagation in irradiated materials
In this section, various modes of crack propagation and defect interaction across different irradiation levels are examined in detail. Section 3.1 examines characteristic modes of crack growth, highlighting how radiation-induced defects alter crack trajectories and lead to phenomena such as crack branching, void coalescence, and secondary crack formation. Section 3.2 delves into the fundamental mechanisms governing defect–crack interactions, focusing on how isolated defects (e.g., voids, dislocation loops, rigid inclusions) either facilitate or impede crack propagation and how these interactions influence fracture behavior. Finally, defect-driven fracture mechanisms are summarized in Section 3.3.